# Memory-Efficient Community Detection on Large Graphs Using Weighted Sketches

### v1

@@ 3.6. Boyer-Moore (BM) majority vote algorithm
The Boyer-Moore (BM) majority vote algorithm (Boyer and Moore, 1991) efficiently identifies the majority element in a sequence — defined as an element that appears more than $n/2$ times in a list of $n$ elements. It was developed by Robert S. Boyer and J. Strother Moore, and published in 1981. The algorithm maintains a candidate for the majority element along with a counter for its ”votes.” Initially, we set the candidate and count variables. As we iterate through the list, if the count is 00, we assign the current element as the new candidate and set the count to $1$. If the count is not zero and the current element matches the candidate, we increment the count; if it differs, we decrement the count. By the end of the iteration, the candidate variable holds the potential majority element. This algorithm is efficient, completing in a single pass with $O(n)$ time complexity and using constant space $O(1)$.
@@ 3.7. Misra-Gries (MG) heavy hitters algorithm
The Misra-Gries (MG) heavy hitters algorithm, introduced in 1982 by Jayadev Misra and David Gries (Misra and Gries, 1982), extends the Boyer-Moore majority finding algorithm for the heavy-hitter problem. It aims to find elements that occur more than $\frac{n}{k+1}$ times, where $n$ is the total number of processed elements and $k+1$ is a user-defined threshold. The algorithm initializes an empty set of up to $k$ counters, each tracking a candidate element and its approximate count. As elements are processed, if an element is already a candidate, its counter is incremented. If it is not a candidate and fewer than $k$ counters are used, a new counter is created with a count of $1$. If all $k$ counters are occupied, the algorithm decrements each counter by $1$. Counters that reach zero result in the corresponding element being removed from the candidate set. After processing the stream, the remaining candidates are those that likely exceed the $\frac{n}{k}$ threshold, although a verification step may be needed to confirm their exact counts. The decrementing mechanism effectively prunes infrequent elements from the list. The MG algorithm operates with a time complexity of $O(n)$ and a space complexity of $O(k)$, making it suitable for scenarios with limited computational resources.
@@ 4. Approach
Our first attempt at reducing the memory usage focuses on scaling down the size of the hashtable’s values array by a factor of 3000×3000$\times$3000 ×. Our findings indicate that this has only a minimal effect on the quality of generated communities. We elaborate on this method in Section A.2. We now move on to present our main method proposed in this paper, which further reduces the memory required for each hashtable to just $1$ KB or less, irrespective of graph size.
@@ 5.1.3. Dataset
Table 1. List of $12$ graphs retrieved from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (with directed graphs indicated by *). Here, $|V|$ denotes the number of vertices, $|E|$ denotes the number of edges (after making the graph undirected by adding reverse edges), and $|\Gamma|$ denotes the number of communities obtained with Static Leiden algorithm (Sahu et al., 2024).
@@ 5.3. Performance of our MG-based Leiden
Next, we assess the performance of the parallel implementation of MG64 Leiden (MG-based Leiden with $k=64$ slots) and compare it to Default Leiden (Sahu, 2023a) and BM-based Leiden, across the graphs listed in Table 1. Figure 8 displays the runtime and the modularity of the communities identified by the algorithms for each graph in the dataset. As shown in Figure 8(a), MG64 Leiden is on average 3.19×3.19$\times$3.19 × slower than Default Leiden. In terms of modularity, Figure 8(b) indicates that MG64 Leiden achieves a modularity similar to that of Default Leiden. BM Leiden exhibits lower modularity compared to MG64 Leiden, and its performance is particularly poor on social networks and the sk-2005 web graph. This may be attributed to the high average degree of these graphs, and the weak community structure typical of social networks. BM Leiden also converges significantly faster on social networks, as shown in Figure 8(a), while obtaining these poor quality communities — but it also converges faster on the asia_osm road network, while obtaining high quality communities. Thus, we recommend employing BM Leiden for road networks and protein k-mer graphs, but not for web graphs or social networks. Similar to MG8 Louvain however, MG64 Leiden can be reliably utilized across all graph types.
@@ 6. Conclusion
In this work, we proposed to replace the per-thread hashtables in these implementations with a weighted version of the Misra-Gries (MG) sketch. Our experiments showed that MG sketches with $8$, $64$, and $8$ slots for Louvain, Leiden, and LPA, respectively, significantly lower memory usage the memory usage of the implementations — while suffering only up to $1\%$ decrease in community quality, but with runtime penalties of 1.48×1.48$\times$1.48 ×, 3.19×3.19$\times$3.19 ×, and 2.11×2.$1$, respectively. Additionally, we presented a weighted Boyer-Moore (BM) variant for Louvain, Leiden, and LPA, which demonstrated good strong performance on specific graph types. We believe that these approaches, while introducing some runtime overhead, are well parallelizable and hold potential to surpass current memory-intensive techniques in performance on devices with a large number of threads.

### v2

@@ 3.6. Boyer-Moore (BM) majority vote algorithm
The Boyer-Moore (BM) majority vote algorithm (Boyer and Moore, 1991) efficiently identifies the majority element in a sequence — defined as an element that appears more than $n/2$ times in a list of $n$ elements. It was developed by Robert S. Boyer and J. Strother Moore, and published in 1981. The algorithm maintains a candidate for the majority element along with a counter for its ”votes.” Initially, it sets the candidate and count variables. As it iterates through the list, if the count is 00, it assigns the current element as the new candidate and sets the count to $1$. If the count is not zero and the current element matches the candidate, it increments the count; if it differs, it decrements the count. By the end of the iteration, the candidate variable holds the potential majority element. This algorithm is efficient, completing in a single pass with $O(n)$ time complexity and using constant space $O(1)$.
@@ 3.7. Misra-Gries (MG) heavy hitters algorithm
The Misra-Gries (MG) heavy hitters algorithm, introduced in 1982 by Jayadev Misra and David Gries (Misra and Gries, 1982), extends the Boyer-Moore majority finding algorithm for the heavy-hitter problem. It aims to find elements that occur more than $\frac{n}{k+1}$ times, where $n$ is the total number of processed elements and $k+1$ is a user-defined threshold. The algorithm initializes an empty set of up to $k$ counters, each tracking a candidate element and its approximate count. As elements are processed, if an element is already a candidate, its counter is incremented. If it is not a candidate and fewer than $k$ counters are used, a new counter is created with a count of $1$. If all $k$ counters are occupied, the algorithm decrements each counter by $1$. Counters that reach zero, result in the corresponding element being removed from the candidate set. After processing the stream, the remaining candidates are those that likely exceed the $\frac{n}{k}$ threshold, although a verification step may be needed to confirm their exact counts. The decrementing mechanism effectively prunes infrequent elements from the list. The MG algorithm operates with a time complexity of $O(n)$ and a space complexity of $O(k)$, making it suitable for scenarios with limited computational resources.
@@ 4. Approach
Our first attempt at reducing the memory usage focuses on scaling down the size of the hashtable’s values array by a factor of 3000×3000$\times$3000 ×. Our findings indicate that this has only a minimal effect on the quality of generated communities. We elaborate on this method in Section A.2. We now move on to present our main method proposed in this paper, which further reduces the memory required for each hashtable to just $0.5$ KB or less, irrespective of graph size.
@@ 5.1.3. Dataset
Table 1. List of $13$ graphs retrieved from the SuiteSparse Matrix Collection (Kolodziej et al., 2019) (with directed graphs indicated by *). Here, $|V|$ denotes the number of vertices, $|E|$ denotes the number of edges (after making the graph undirected by adding reverse edges), and $|\Gamma|$ denotes the number of communities obtained with Static Leiden algorithm (Sahu et al., 2024).
@@ 5.3. Performance of our MG-based Leiden
Next, we assess the performance of the parallel implementation of MG64 Leiden (MG-based Leiden with $k=64$ slots) and compare it to Default Leiden (Sahu, 2023a) and BM-based Leiden, across the graphs listed in Table 1. Figure 8 displays the runtime and the modularity of the communities identified by the algorithms for each graph in the dataset. As shown in Figure 8(a), MG64 Leiden is on average 3.19×3.19$\times$3.19 × slower than Default Leiden. In terms of modularity, Figure 8(b) indicates that MG64 Leiden achieves a modularity similar to that of Default Leiden. BM Leiden exhibits lower modularity compared to MG64 Leiden, and its performance is particularly poor on social networks and the sk-2005 web graph. This may be attributed to the high average degree of these graphs, and the weak community structure typical of social networks. BM Leiden also converges significantly faster on social networks, as shown in Figure 8(a), while obtaining these poor quality communities — but it also converges faster on the asia_osm road network, while obtaining high quality communities. Thus, we recommend employing BM Leiden for road networks and protein k-mer graphs only. However, MG64 Leiden can be reliably utilized across all graph types.
@@ 6. Conclusion
In this work, we proposed to replace the per-thread hashtables in these implementations with a weighted version of the Misra-Gries (MG) sketch. Our experiments showed that MG sketches with $8$, $64$, and $8$ slots for Louvain, Leiden, and LPA, respectively, significantly lower memory usage the memory usage of the implementations — while suffering only up to $1\%$ decrease in community quality, but with runtime penalties of 1.48×1.48$\times$1.48 ×, 3.19×3.19$\times$3.19 ×, and 2.11×2.$1$, respectively. Additionally, we presented a weighted Boyer-Moore (BM) variant for Louvain, Leiden, and LPA, which demonstrated good performance on specific graph types. We believe that these approaches, while introducing some runtime overhead, are well parallelizable and hold potential to surpass current memory-intensive techniques in performance on devices with a large number of threads.